1.After paying $1.50, you are allowed to open a newspaper vending machine freely (i.e. get as many as you want). In contrast, a soft drink vending machine only drops one can after you pay the same amount of money. Why? Please explain the difference in your own words. use your own words.

2 Assume that a salesman learned the price elasticity of demand for his products is −2.0. How many percents will increase in the total sales (revenue) if he cuts the price by 10%? Show your calculation process. use your own words.
Answer #1

1) We are allowed to take as many newspapers as we want from the vending machine but only one can of soft drink at the same price. This is because of the benefit obtained by taking  two or more newspapers would be very less or nothing to the person reading the newspaper as he wouldn’t be interested in reading the same news from two or more different copies of the same newspaper. In case of soft drink a person would get more utility by consuming more than one unit of soft drink at the price of one unit and will tend to have more and more units as his utility is increasing by having more cans of soft drinks. So, we are allowed to take only one can of soda for $1.50 but there is no limit on the copies of newspaper that we can take for $1.50 from the vending machine as the utility obtained from by consuming more units of soft drink will increase whereas the total utility from taking further copies of newspapers will remain constant.

2) Price elasticity of demand is the responsiveness of the quantity demand to the change in the price of the good. The price elasticity of demand =

Percentage Change in quantity demanded / Percentage change in price

-2 = Percentage change in quantity demanded/ -10

Percentage change in quantity demanded = 20%.

So we can say that due to a cut in prices by 10%, the quantity demanded will rise by 20%.

So, suppose if initially 100 units were sold at the price of $100. The total revenue was 100*100 = $10000.

Now If price decreases by 10% and quantity increases by 20%, new price will be $90 and the new quantity will be 120. So, total revenue = 120*90 = $10800

So Percentage change in revenue = (Change in revenue/ Initial revenue) *100

= ((10800-10000)/10000) * 100

= 8%.

So we can say that the revenue will increase by 8%.

IF YOU FIND THIS ANSWER HELPFUL, PLEASE UP VOTE.